state the value of the discriminant and number of real solution s^+3s+3=0 and one last? the cost of maufactor amolded part is related to the quanity produced during a production run. when 100 parts are produced, the cost is 300. when 500 parts are produced, the cost is 1900 dollars wat is the averae cost per part.
Answer by drk(1908)

state the value of the discriminant and number of real solution s^+3s+3=0
discriminant is 9-12 = -3
There are zero real solutions.
---
the cost of maufactor a molded part is related to the quanity produced during a production run. when 100 parts are produced, the cost is 300. when 500 parts are produced, the cost is 1900 dollars wat is the averae cost per part.
Our coordinates are (100, 300) and (500, 1900)
average = (1900-300)/(500-100) = 1600/400 = $400/ part